OpenStudy (texaschic101):

I am having trouble with this....I have done it twice and am getting different answers...help

OpenStudy (texaschic101):

If 3x + (1/3)x = 3.....find the value of 9x^2 + 1/9x^2

OpenStudy (texaschic101):

oops...the last part of that is (9x^2 + 1)/(9x^2)

jimthompson5910 (jim_thompson5910):

what do you get when you solve 3x + (1/3)x = 3 for x

OpenStudy (texaschic101):

I got 9/4 = x

jimthompson5910 (jim_thompson5910):

that's not correct

OpenStudy (texaschic101):

let me do it again... 3x + 1/3x = 3 3/3x + 1/3x = 3 4/3x = 3 x = 3 * 3/4 = 9/4 what am I doing wrong ?

jimthompson5910 (jim_thompson5910):

3x is NOT 3/3x

jimthompson5910 (jim_thompson5910):

3x is (9/3)x

OpenStudy (texaschic101):

oh....9/3x... 9/3x + 1/3x = 3 10/3x = 3 x = 3*3/10 = 9/10

OpenStudy (texaschic101):

should I make that a decimal for the other problem or leave it as a fraction ?

jimthompson5910 (jim_thompson5910):

leave it as a fraction

jimthompson5910 (jim_thompson5910):

now plug it into (9x^2 + 1)/(9x^2) and simplify

OpenStudy (texaschic101):

(9x^2 + 1)/(9x^2) (9(9/10*9/10) + 1)/ (9(9/10 *9/10) right so far ?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you replaced every x with 9/10 and expanded (9/10)^2 into (9/10*9/10)

OpenStudy (texaschic101):

(9(81/100) + 1)/(9(81/100) ((729/100) + 1) / (729/100) (729/100 + 100/100) / (729/100) (829/100) / (729/100) 829/100 * 100/729 829/729......is that correct ?

jimthompson5910 (jim_thompson5910):

very nice work

OpenStudy (texaschic101):

It all started with that stupid mess up at the very beginning.....thanks for setting me straight

jimthompson5910 (jim_thompson5910):

yw, glad to be of help

OpenStudy (texaschic101):

I will close this now.....jim, your the best :)

jimthompson5910 (jim_thompson5910):

thanks lol